### v7

# Interpreting GNN-based IDS Detections Using Provenance Graph Structural Features

Abstract
Advanced cyber threats (e.g., Fileless Malware and Advanced Persistent Threat (APT)) have driven the adoption of provenance-based security solutions. These solutions employ ML models for behavioral modeling and critical security tasks such as malware and anomaly detection. However, the opacity of ML-based security models limits their broader adoption, as the lack of transparency in their decision-making processes restricts explainability and verifiability. We tailored our solution towards Graph Neural Network (GNN)-based security solutions since recent studies employ GNNs to comprehensively digest system provenance graphs for security critical tasks.
To enhance the explainability of GNN-based security models, we introduce ProvExplainer, a framework offering instance-level security-aware explanations using an interpretable surrogate model. ProvExplainer’s interpretable feature space consists of discriminant subgraph patterns and graph structural features which can be directly mapped to the system provenance problem space, making the explanations human understandable. Considering both the subgraph patterns and graph structural features, gives ProvExplainer the unique advantage of providing explanations that are sensitive to both local and global contexts.
By considering prominent GNN architectures (e.g., GAT and GraphSAGE ) for anomaly detection tasks, we show how ProvExplainer outperformed the current State-Of-The-Art (SOTA) GNN explainers to deliver security domain and instance-specific explanations. We measure the explanation quality using $\text{fidelity}^{+}$ / $\text{fidelity}^{-}$ metric as used by traditional GNN explanation literature, we incorporate the precision/recall metric where we consider the accuracy of the explanation against the ground truth, and we designed a human actionability metric based on graph traversal distance. On real-world Fileless Malware and APT datasets, ProvExplainer achieves up to 29%/27%/25%/1.4x higher $\text{fidelity}^{+}$ , precision, recall and actionability (where higher values are better), and 12% lower $\text{fidelity}^{-}$ (where lower values are better) when compared against SOTA GNN explainers.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2024 26 February - 1 March 2024, San Diego, CA, USA ISBN 1-891562-93-2 https://dx.doi.org/10.14722/ndss.2024.23xxx www.ndss-symposium.org
I Introduction
Given the significant threat posed by advanced and sophisticated adversaries [1, 2, 3, 4] (i.e., malware writers and APT actors), security is paramount in today’s highly digitized society. Recent advances in system monitoring have generated a variety of fine-grained telemetry for advanced security analysis. To leverage this telemetry data to counter threats, various learning-based security tools [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] have been deployed, with provenance-based solutions being the most notable.
System provenance captures causally dependent system events on an host, represented as a provenance graph. Provenance graphs are heterogeneous, comprising various node types (e.g., process, file, and socket) and edge types (e.g., read, write, and send), each with textual and numerical attributes, effectively representing dynamic runtime behaviors. Hence, provenance graphs provide irreplaceable security defense for countering sophisticated and stealthy APT attack campaigns. Particularly, Endpoint Detection and Response (EDR) solutions [22, 23, 24], built on fine-grained system provenance data, have become a mainstream security defense for enterprises.
Recent advancements in GNNs enable the direct learning of relational dependencies and topological structures from graphs. GNNs utilize a permutation-invariant message-passing mechanism to aggregate information from node neighborhoods, expanding the information propagated through each neighborhood with every iteration. This allows GNNs to learn from complex graph structures automatically, which has proven beneficial for provenance solutions [21, 20, 19, 25, 26, 27].
While GNN-based security models are highly effective, their opaque nature significantly limits their advantages, i.e., due to the lack of clear and verifiable explanations. These models fail to provide ground truth relevant explanations, which are essential for building trust among security practitioners. In this context, ground truth relevant explanations refer to explanations aligned with specific system actions that represent Tactics, techniques, and Proceduress (TTPs), and system artifacts that are familiar to analyst and can be validated by existing security intelligence (e.g., attack memos [28, 29] and VirusTotal [30] reports).
While studies [31, 32, 33] have explored domain-agnostic explanations for GNN decisions, they lack validation in the security domain, where contextualizing explanations with program behaviors is critical for trusting the underlying model. Recent research [34, 35, 36] has emphasized verifiable explanations for neural networks in binary analysis and vulnerability discovery, yet our study is the first to validate explanations within the system provenance domain.
Explanations have different meanings for model developers and end-users, since end-users cannot make sense of explanations in terms of model weights, but can make sense if explanations are in terms of edges or subgraphs. Therefore, in the security domain, models must explain their decision-making processes [37] that are verifiable using the ground truth and domain expertise.
In this paper, we introduce ProvExplainer, a novel approach to enhance the explainability of GNN-based detection models trained on system provenance data by utilizing an interpretable surrogate model, specifically Decision Trees (DT), with security-aware discriminant subgraph patterns and graph structural features. We efficiently mined subgraph patterns [38] using a graph evolution miner (GERM [39]) to identify discriminative subgraph patterns between malicious and benign datasets. Our analysis revealed that these subgraph patterns correspond to well-known malicious activities, such as system probing [40], and malware replication [41] and staging [42]. Additionally, we extracted structural graph features, such as degree centrality, betweenness centrality, closeness centrality, eigenvector centrality, clustering coefficient, and clustering triangles. These structural features capture long-term dependencies within the system, while the subgraph patterns represent local dependencies. Together, they provide explanations for both short-term and long-term dependencies.
ProvExplainer is specifically designed to address the unique challenges posed by provenance graphs—complex, heterogeneous, and large-scale structures that capture dynamic system behaviors with fine-grained causal semantics. Unlike prior efforts that apply explanation techniques to general-purpose graph domains, our work introduces a domain-specialized explanation framework tailored for provenance-based GNN intrusion detection. Provenance graphs differ significantly from other graph modalities due to their temporal evolution, multi-typed nodes and edges, and security-relevant semantics embedded in system events. These aspects introduce non-trivial complexities, such as the need to capture temporal causality, reason over multi-hop dependencies, and interpret system-level semantics in a trustworthy manner.
ProvExplainer addresses these challenges by adapting DT-based surrogate models with domain informed features, discriminative subgraph patterns and graph-theoretic measures, that are aligned with operational security workflows. For security analysts, ProvExplainer offers three critical benefits: (1) it builds confidence that the GNN bases its predictions on legitimate anomalous subgraphs in the provenance graph; (2) it enables analysts to investigate alerts by localizing explanations to meaningful subgraphs representative; and (3) it equips operators with interpretable justifications to support triage and response decisions.
We emphasize that ProvExplainer is an explanation framework, not an attack detection technique. Unlike GNNs, surrogate Decision Trees (DTs) sacrifice generalizability, expressiveness, and automatic feature extraction to gain interpretability, making them unsuitable for direct attack detection. ProvExplainer is a GNN-structure-agnostic framework capable of incorporating an arbitrary number of subgraph patterns and structural graph features, and it is designed for easy extensibility with additional security-aware graph features.
We focus on the structural features of system provenance graphs, without considering the text or numerical node and edge attributes. We restrict the scope of our attribute view for the following reasons: first, the core merit of provenance-based Intrusion Detection System (IDS) lies in the ability to retain features drawn from causal dependencies captured in graph structural relations; these structural features are significantly more difficult for adversaries to manipulate than textual attributes (e.g., filenames). Secondly, if we cannot explain GNN decisions based solely on graph structures, it becomes even more challenging to interpret decisions when GNNs incorporate both structural and attribute-based complexities.
We evaluated the effectiveness of ProvExplainer using $\text{fidelity}^{+}$ $\text{fidelity}^{-}$ , following prior GNN explainability research [43]. An effective explainer should have a higher $\text{fidelity}^{+}$ , as it quantifies the prediction change when important subgraphs are removed, and a lower $\text{fidelity}^{-}$ , as it reflects the prediction change when non-important subgraphs are excluded. We also incorporated two additional measures, precision and recall, that captures the accuracy of the explanation provided compared against the ground truth. Finally, we introduce a human actionability metric that measures the graph traversal distance between an explanation and the ground truth artifacts; a lower score indicates that the explanation guides analysts more directly to the relevant evidence, thereby reducing their investigative workload. In our experimental evaluation, we show ProvExplainer surpasses current SOTA GNN explainers in the majority of our security-oriented evaluation tasks.
To demonstrate the advantages of ProvExplainer in the system provenance domain, we compared ProvExplainer against SOTA GNN explainers such as GNNExplainer [31], PGExplainer [32], and SubgraphX [33] across well-balanced security domain tasks, including malware detection, and APT detection. These SOTA GNN explainers provide the best performance according to other security studies [44, 35]. The GNN models in our evaluation are trained with an extensive dataset collected using our in-house deployment which monitored system events on an average of 86 mixed Windows and Linux hosts over 13 months. In addition, in our evaluation, we used the DARPA [darpa] Transparent Computing dataset, the APT dataset from [46], and malware lists from [47].
To the best of our knowledge, our research is the first to leverage security-aware discriminant subgraph pattern and graph structural features specifically tailored for human interpretation of GNN-based detection models. We summarize the contributions of ProvExplainer as follows:
  • ProvExplainer utilizes discriminant subgraph patterns and graph structural features to provide explanation for GNN-based detection models.
  • Our security-aware features enabled surrogate DTs to achieve 95% agreement with the GNN detection model.
  • ProvExplainer increased $\text{fidelity}^{+}$ (e.g., higher is better) by 29%, precision by 27%, recall by 25%, human actionability by 1.4x and decreased $\text{fidelity}^{-}$ (e.g., lower is better) by 12% compared to SOTA GNN explainers.
  • We curated an extensive dataset of APT scenarios ( i.e., DARPA [darpa] and past literature [46]) and collected real-world malware samples from [47] to confirm the generalizability of ProvExplainer.
II Background and Related Work
II-A System Provenance and Data Collection
System provenance analysis [king:2003sosp, 49] leverages data collection agents on end-hosts to collect interaction events among key system resources: processes, files, and network sockets. This work relies on in-house data from 86 hosts in a university environment, the DARPA Transparent Computing dataset [darpa], and datasets from previous study [46]. Our in-house data collection accumulates 13 GB to 92 GB daily, tracking around 875 unique programs, 7,025K processes, 4,824K network connections, and 111,583K file operations.
Our system provenance data schema, detailed in Table III, is similar to DARPA’s Common Data Model (CDM) [darpa] schema, but we omit memory objects, registry events, and thread distinctions within a process. These choices were made to balance real-world overhead constraints of load balancing and storage. We also established a malicious testbed to generate malware execution traces. To ensure the freshness and realism of our malware samples, we utilize Cyber Threat Intelligence (CTI) feeds [attck], the VirusTotal database [vtotal], and penetration testing TTP [metasploit, cobaltstrike, immunity]. We refer to DARPA TC dataset [darpa] and previous literature of [46] for APTs.
II-B Provenance-based ML Security Solutions
Recent advancements in system event collection have revitalized host-based IDSs. Although initially proposed in the late ’90s [54], host-based IDSs have proved effective against advanced attacks such as APTs and Fileless Malware . Various studies [15, 18, 55, 56, 19, 20, 21, 26, 27, 57] demonstrate the efficacy of provenance-based Machine Learning (ML) techniques in identifying behavioral deviations. However, the high detection performance is offset by a lack of trustworthy explanations and high false positive rates, impacting operational confidence and resource allocation. Our work focuses on enhancing the transparency and explainability of provenance-based ML security. We concentrate on a balanced set of security tasks of malware and APT detection. In this paper, we consider graph-level tasks; providing explanations for node and edge level tasks is a promising direction for future work that aligns with upcoming IDS trends.
GNNs for System Provenance. Our research employs an industry-standard GNN framework (i.e., Deep Graph Learning (DGL)) [58] to model and explain system provenance graphs. Leveraging DGL’s mature development ecosystem ProvExplainer aligns with current analytical techniques and streamlined integration into real-world applications. Despite the security community’s historical preference for custom detectors [55, 56, 19, 20, 21, 25, 27, 26] due to the complexity and heterogeneity of provenance graphs, we chose a general GNN framework for universal applicability and broader impact. Notably, the DGL community has integrated our heterogeneous GNN enhancement.
II-C Explainability and Security Applications
Explaining ML-based Security Models. Due to the importance of explainability in the security domain, several explainability frameworks have been proposed for ML-based security solutions. LEMNA [guo2018lemna] focused on classifying PDF malware and detecting a function’s entry point in binary code using regression mixture models as a localized surrogate to approximate the classifier’s decision boundary. CFGExplainer [60] use deep surrogate models to explain GNNs but work exclusively work on homogeneous graphs.
Jacob et al. [jacobs2022ai] proposed TRUSTEE, a framework that generates decision tree (DT)-based interpretations for ML models to detect shortcut learning. Their success in clarifying model decisions about network packets inspired us to generalize the approach to the system provenance domain. System provenance domain relies on heterogeneous multi-attributed graph datasets, which are not natively consumed by DTs.
Explaining GNN Models. Recent research in GNN explainers [31, 32, 33] has advanced in identifying key nodes, edges, or subgraphs in GNNs, and are categorized into white-box and black-box explainers. White-box methods, e.g., GNNExplainer [31] and PGExplainer [32], access GNN internals, including model weights and gradients. Conversely, black-box methods like SubgraphX [33] operate on model inputs and outputs, reducing coupling between the explanation framework and model architecture. [62] has shown high variance in explanations provided by traditional GNN explainers, raising reliability and applicability concerns.
II-D Ground-truth Verification
Ground truth verification of ML model decisions in security-critical tasks [34, 35, 36] has garnered significant attention, highlighting the role of verifying explanations in unveiling the model’s decision-making process. These studies highlight the effectiveness of white-box techniques in malware and vulnerability discovery, and reconstructing ground truth around local explanations [36].
In system provenance, “ground truth” refers to the real-world system information (e.g., system actions) against which the validity of a model’s predictions is checked. In this paper, we approximate the ground truth using documentation created by security vendors, tech reports, and previous studies. The relevant processes, files, and network sockets and their associated events mentioned in the documentation are designated as documented entities.
We extract documented entities with three methods: (1) we refer to pre-existing malware profile databases that contain information from different security vendors, such as VirusTotal [30], to obtain activity summaries detailing network communications, file system actions, and process behaviors; (2) we extract key entities (i.e., process involved, files created and connections made) from tech reports [28, 29] and system manuals [63]; (3) we consult dataset authors [46] and review dataset documentation to identify the components of each attack present in the datasets. We understand that the documentation provided by the security vendors can have experimental errors, as well as selection and experimental bias. To mitigate these problems, we aggregated information from multiple sources.
III Motivation
In this section, we present a real-world attack scenario from the DAPRA TC [29] dataset to highlight the limitations of existing GNN explainers and demonstrate the effectiveness and utility of ProvExplainer. The experimental setup to evaluate ProvExplainer is described in § VI and more case studies are described in § VII and appendix (§ A-E and § A-F).
(a) Attack Graph.
(b) SOTA GNN explainations.
(c) Pattern #1 and Pattern #2.
Figure 1: Trace: after an employee clicks on a phishing link, Firefox installs multiple Trojans to exfiltrate sensitive data.
Attack Scenario (Trace: Phishing Email [darpa]). Figure 1 presents a simplified depiction of the APT attack initialized by phishing e-mail link. This particular attack was executed as part of a DAPRA red team exercise by Trace team. First, the attacker launches a phishing campaign that compromises the identity of an employee. Leveraging this compromised identity, the attacker then sends deceptive emails containing links to a malicious website (i.e., www.nasa.ng) to other employees. When an employee visits the website, a Trojan is downloaded and placed in /tmp/vUgefal. Upon execution, the Trojan replicates itself within the system by reading from its own file.
The malware adopts the name firefox nasa.ng and stores sensitive content in /usr/share/locale/locale.alias for exfiltration. To evade detection, the Trojan creates additional malicious copies using benign program names (e.g., firefox), allowing it to replicate unhindered and overwhelm the system with malicious processes. These malware clones quietly read sensitive user files while the original Trojan remains undetected.
Limitation of SOTA GNN explainers. 1(b) compares the explanations produced by ProvExplainer with those of SOTA GNN explainers (i.e., GNNExplainer, PGExplainer, and SubgraphX) in capturing malicious activities relevant to GNN-based IDS. While GNNExplainer effectively uncovers the initial compromise behavior such as payload dropping and staging, it fails to capture later attack stages notable the interactions with shared libraries and cache files. PGExplainer, in contrast, highlights irrelevant benign system components like xfce4-appfinder (invoked by a DARPA script to simulate an enterprise environment) but overlooks critical malicious interactions. SubgraphX accurately detects the malicious inheritance behavior of Firefox creating multiple malware clones, yet it misses key points such as how the malware writes its payload to disk and accesses sensitive files.
ProvExplainer presents a complete view of the kill chain from Trojan deployment and clone creation to the final step of accessing sensitive data. ProvExplainer uses security relevant discriminative subgraph patterns and graph structural metrics as features to capture the holistic malicious behavior within the system. Moreover, it accomplishes this despite functioning as a black-box solution, requiring neither direct access to the underlying model nor the training data. This flexibility makes ProvExplainer compatible with a wide range of GNN-based IDS deployments.
ProvExplainer Features: Clustering coefficient (graph structural metric); pattern #1 and #2 (discriminative subgraph patterns).
System Interpretation of ProvExplainer Features. In this APT scenario, clustering coefficient, pattern #1 and pattern #2 in 1(c) showcase the malware’s structures. Pattern #1 reveals the Trojan’s propagating activity, where initial malware (Process P1) reads the firefox binary (/usr/local/bin/firefox) to execute it and write the maliclious payload (/tmp/vUgefal) to a file that is read by other malicious firefox processes (Process P4, P7,P8, P9). A series of intra-malware communication happens when the another malicious firefox (Process P2) reads the malicious payload and writes to system file ( /etc/nsswitch.conf). As seen in pattern #2, other malware processes (Process P4, P5, P6) read the system file to update their configurations and task lists to better attack the victim. This propagation of information continues such that any subsequent malware processes that read from the system file gets the updated configuration.
By masquerading as benign programs, the malware disguises its malicious processes, allowing it to proliferate undetected. The clustering coefficient of the graph increases because of multiple malware reading the template files by forming large process creation and file read clusters.
IV Problem Statement and Threat Model
Our research addresses explainability in GNN-based security models built on system provenance graphs, tackling a core issue in the security domain. The complexity of explaining GNN decisions is exacerbated by graph structural learning, which adds to the inherent complexity of Neural Networks (NNs). Existing studies on GNN explainability [43, gnnexplainer, subgraphx, pgexplainer, 60, 36] often fail to effectively map back to system behaviors in the provenance domain. To bridge this gap, ProvExplainer, employs a surrogate DT-based method that utilizes interpretable security-aware graph structural features (i.e., discriminant subgraph patterns and graph structural features) to explain GNN decisions. Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and a GNN model $\mathcal{M}$, an explanation method $\mathcal{EM}$ produces:
  • Set of subgraphs $\mathcal{SG}$, where $\mathcal{SG}=\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{k}\}$ and $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$ for $i=1..k$, with $\mathcal{V}_{i}\subseteq\mathcal{V}$ and $\mathcal{E}_{i}\subseteq\mathcal{E}$.
  • Set of properties $\mathcal{P}rop=\{p_{1},p_{2},\dots,p_{j}\}$, where each $p_{j}$ is a graph structural property of $\mathcal{G}$.
The subgraphs $\mathcal{SG}$ and properties $\mathcal{P}rop$ are designed to highlight the most relevant components of $\mathcal{G}$ that contribute to the predictions of $\mathcal{M}$.
Our threat model assumes the integrity of on-device data collection, relying on provenance records secured by existing mechanisms ([provdetector, 18, nodoze, 46, 65, 19, 20, 21]). Our primary objective is to generate security-aware explanations to aid security analyst and increase their trust in the GNN’s decisions. We consider graph-level anomaly detection tasks; explaining GNN decisions in node/edge level tasks is outside the scope of this work. Systematically aggregating an accurate and trustworthy ground truth for malware and APT behaviors poses a challenging open problem. In this paper, we approximate the ground truth using publicly available documentation (§ II-D). In line with recent literature on GNN explanation [60, 34, 35], adversarial samples are outside the scope of the paper. Creating robust detection and explanation systems that can withstand adversarial manipulation [goyal2023sometimes, 46], dataset poisoning, and detection model manipulations are critical open research problems that are orthogonal to our work.
V ProvExplainer
Figure 2: ProvExplainer architecture.
We apply ProvExplainer in three stages (Figure 2).
Stage 1: Extract Security-aware Features (§ V-A). We extract security-relevant features such as subgraph patterns using a discriminative subgraph miner [38] and graph structural features such as clustering and centrality measures that facilitate distinction between benign and anomalous datasets. The subgraph patterns enable the identification of localized attack vectors, while global graph structural features capture holistic attack behaviors.
Stage 2: Train an Interpretable Surrogate Model (§ V-B). Next, we utilize diverse system provenance dataset to train an interpretable surrogate DT that aligns with the GNN decisions using the extracted features. Since the GNN’s decision-making process relies on a structure-dependent message-passing algorithm, we efficiently mimic the decision-making process of the GNN’s using the security-relevant features.
Stage 3: Interpret the Surrogate Model (V-C). To interpret the explanation for detection using the surrogate DT, we extract the graph nodes participating in the subgraph patterns that contribute to the DT’s decision. These explanations are valid only when the surrogate agrees with the GNN.
V-A Feature Extraction
Algorithm 1 Discriminative Pattern Mining
1:Benign $D_{b}$ and Anomaly $D_{a}$ datasets, $k$, and score_funcscore\_func
2:Top-$k$ Discriminative Subgraph Patterns
[…]
5:for each pattern $p\in CP$ do
6:  $div\leftarrow\text{getDIVERSITY}(p)$
7:  for each graph $G\in D_{b}$ do
8:    $size_{b},support_{b}\leftarrow\text{getSIZEandSUPPORT}(p,G)$
9:  end for
10:  $cov_{b}\leftarrow\text{getCOVERAGE}(p,D_{b})$
11:  $score_{b}\leftarrow score\_func(div,cov_{b},size_{b},support_{b})$
12:  for each graph $G\in D_{a}$ do
13:    $size_{a},support_{a}\leftarrow\text{getSIZEandSUPPORT}(p,G)$
14:  end for
15:  $cov_{a}\leftarrow\text{getCOVERAGE}(p,D_{a})$
16:  $score_{a}\leftarrow score\_func(div,cov_{a},size_{a},support_{a})$
17:  $Patterns.\text{append}([p,|\text{score}_{a}-\text{score}_{b}|])$
18:end for
19:Sort PatternsPatterns in desc order based on the score difference
20:return $Patterns[:k]$
Security-relevant Features. The extraction of security-relevant subgraph features is pivotal in distinguishing anomalous patterns from benign patterns within provenance datasets. To avoid information leakage and ensure no data snooping, patterns that are mined exclusively on the training dataset are used for explanation evaluation. This approach enables unbiased pattern discovery and allows the identified discriminative features to be evaluated solely on the testing dataset, maintaining a rigorous evaluation protocol.
Frequent subgraph patterns are mined from the training dataset using a graph evolution miner (GERM [39]), where subgraphs of specified sizes and their supports are identified. Support means how many times the subgraph instances were identified in the graph. These patterns serve as representative features, capturing the frequency and structure of motifs present in the datasets.
For each subgraph pattern, we calculate its pattern diversity, which measures the structural complexity of the pattern. Diversity scores are computed by aggregating node and edge weights, where node weights are assigned according to unique node labels, and edge weights reflect the types of connections between nodes. This diversity score quantifies the variability within each pattern, aiding in highlighting complex, potentially discriminative structures.
We also measure the coverage of a subgraph pattern by measuring its presence across graphs within the dataset, providing a statistical representation of how commonly the pattern occurs in the dataset as a whole. Calculated as the proportion of graphs in which the pattern appears with non-zero support, coverage serves as an indicator of the pattern’s prevalence or rarity in benign and anomalous instances. High coverage patterns suggest commonly occurring structures within the respective dataset, representing anomalous or benign operations. We want to identify patterns that have high coverage in anomalous but low coverage in benign datasets, providing a basis for distinguishing between them.
To quantify the discriminative utility of each pattern, we define a scoring function parameterized by weights $\alpha,\beta,\gamma,\text{ and }\eta$, which represent the relative importance of diversity, coverage, pattern size, and support, respectively. The score_func in Algorithm 1 (i.e., scoring function) is:
$\text{Score}=\alpha\cdot\texttt{div}+\beta\cdot\texttt{cov}+\gamma\cdot\texttt{size}+\eta\cdot\texttt{support}$
where div denotes the diversity of the pattern, cov represents the proportion of graphs in which the pattern appears, size is the average number of nodes in the pattern, and support reflects the frequency of pattern occurrence within each graph.
Increasing $\alpha$ enhances the selection of patterns with diverse node types, which may improve the ability to identify unique patterns but could reduce the focus on more frequent patterns. Raising $\beta$ results in higher scores for patterns that appear across a larger number of graphs, promoting generalization but potentially overlooking rare but important patterns. Elevating $\gamma$ favors patterns with larger node counts, which can capture more complex interactions but might exclude simpler yet crucial substructures. Finally, increasing $\eta$ prioritizes patterns with higher frequency, amplifying commonly occurring patterns but at the expense of pattern diversity.
For our study, we used the following weights: $\alpha=0.2$, $\beta=0.4$, $\gamma=0.1$, and $\eta=0.3$. These values were selected empirically to prioritize coverage and support as the primary differentiators, while also ensuring that the extracted patterns are diverse and large. Subgraph patterns are extracted from each graph in the dataset using Algorithm 1 but the discriminant subgraph patterns are selected exclusively from the anomalous dataset, as we ensure that patterns contributing significantly to the anomalous dataset’s structure are prioritized. Patterns that maximize the score difference between benign and anomaly datasets are selected as discriminative patterns, as they effectively differentiate between benign and anomalous behaviors. Also, subgraph patterns are extracted solely from the training set and evaluated on an independent test set, thereby eliminating any risk of data leakage.
Graph Structural Features. From graph theory literature, we selected the following structural features [67]: degree centrality, closeness centrality, betweenness centrality, eigenvector centrality, clustering coefficient, and clustering triangles. These standard features are known for capturing the key characteristics of graphs, serving as a custom signature that effectively reflects the global properties of a graph. To refresh the reader, we provide informal definitions and their meaning in system provenance domain.
Degree Centrality. Degree centrality is simply the degree of the node, or the count of its incident edges. The degree centrality of a node indicates the extent of its direct connections in the graph. When aggregated across all nodes, high degree centrality values suggest a densely connected system with significant interactivity, whereas low values indicate sparsely connected regions, potentially reflecting isolated or specialized nodes. High degree centrality could suggest the presence of core process hubs, while low values imply a more dispersed system with no central process hubs. However, degree centrality alone cannot definitively indicate malicious behavior. For example, a long-running attack creating a long process chain shows a low degree of centrality, while attacks that interact with many files or fork multiple child malware processes will exhibit a high degree of centrality. In the case of a cryptominer masquerading as schtasks.exe, we expect the benign schtasks.exe to show low centrality as it schedules tasks, while the cryptominer malware have high centrality due to spawning multiple malicious processes.
Closeness Centrality. Closeness centrality is the inverse of the mean shortest path distance from all other nodes to the target node, indicating how close the nodes are within the graph. Closeness centrality measures how centrally located a node is within the graph by examining its proximity to other nodes. High closeness centrality across the graph suggests that many nodes are relatively close to each other, indicating efficient communication pathways. In contrast, low closeness centrality values highlight dispersed sections of the graph, potentially representing decentralized or peripheral interactions within the system. High closeness centrality indicates well-connected process hubs central to the system, while low values suggest that processes are widely dispersed with limited central hubs. Empirically, system processes often have high closeness centrality, while user-level or malware processes show lower values. For instance, botnets masquerading as system-level processes will be distinguishable, since botnet will have a higher closeness centrality due to multiple network connecting being created due to their DDoS activities.
Betweenness Centrality. Betweenness centrality measures the frequency with which a node appears on the shortest paths between other nodes, highlighting its role as a waypoint. It indicates how often a node acts as a bridge along the shortest paths between other nodes. High betweenness centrality suggests a system’s reliance on certain core hubs or bottleneck paths, while lower values indicate multiple alternative pathways and a more distributed structure. High betweenness centrality indicate processes that serve as intermediaries in data or control flow, whereas low values suggest a network of distributed pathways with minimal reliance on specific process hubs. Malware like worms and spyware tend to have low betweenness centrality, as they seek to spread across the system and create multiple distributed information pathways.
Eigenvector Centrality. Eigenvector centrality assigns higher values to nodes connected to other influential nodes, reflecting the prominence of node clusters that are influential. High eigenvector centrality in processes indicates hubs with extensive control over information flow, while low values imply a flatter structure without prominent hubs. Ransomware have high eigenvector centrality, as malicious processes involved in reading and encrypting files act as critical nodes. They keep few child processes to allow high-volume file encryption without overloading the system.
Clustering Triangles. Clustering triangles count the number of triangles that include the node. The clustering triangles count represents the frequency of triangular connections in the graph. High clustering triangle values indicate localized connectivity, suggesting shared resources or close interactions among nodes, while low values indicate a more hierarchical or linear flow. Malware often creates prominent triangles when they write template files and repeatedly execute them, making copies of themselves. Similarly, legitimate programs like svchost.exe (a service scheduler) form triangles when creating conhost.exe and reading standard *.dll files. In contrast, ransomware reads files to encrypt them, creating “spoke-and-wheel” shape rather than triangles, resulting in lower triangle counts than benign processes, svchost.exe.
Clustering Coefficient. The clustering coefficient is calculated by dividing the number of triangles the node participates in by its degree, indicating how close its connections form a complete graph with its neighbors. The clustering coefficient measures the degree to which nodes’ neighbors are interconnected. A high clustering coefficient suggests that nodes form close-knit groups, often indicating shared resources or community-like interactions, while a low coefficient reflects more dispersed connections. High clustering coefficients point to strong groupings around key processes with shared resources, while low values suggest processes with less close-knit hubs. The clustering coefficient provides a similar interpretation to closeness centrality; however, while closeness centrality highlights a process’s accessibility within the system, the clustering coefficient emphasizes the density of its immediate neighborhood connections.
V-B Surrogate Model Training
To obtain explanations from our security-aware graph structural features, we use them to train a global surrogate DT. By training a DT to agree with the predictions of a GNN model, we can interpret the DT to gain insights about the GNN’s decision-making process. To achieve the best agreement results, we enhance traditional DT training with data augmentation that iteratively increases the weight of incorrectly classified samples [61].
We begin with a labeled set of graphs $D_{G}=(G,Y)$, which is used to train the GNN. GNN’s predictions on $D_{G}$ are collected, yielding $GNN(D_{G})=Y^{\prime}$. To prepare the dataset for training the DT, we extract the discriminative subgraph patterns and graph structural features (§ V-A) $F$ and associate them with the GNN’s predictions $Y^{\prime}$ to create a labelled feature dataset $D_{F}=(F,Y^{\prime})$, which we split into train, validation, and testing to evaluate the surrogate DT.
V-C Surrogate Model Interpretation
We obtain instance-level explanations after training the surrogate DT. Each decision node in the DT corresponds either to a discriminant subgraph pattern or to a graph-structural measure (with its associated threshold). Nodes closer to the root of the tree exert greater influence on the decision path; therefore, we rank subgraph-participating nodes based on their relative position in the DT.
In cases where (1) the surrogate model’s prediction diverges from that of the GNN, and (2) the decision path is composed solely of structural measures without subgraph patterns, node-level explanations cannot be extracted, even though a DT is generated. While structural metrics alone cannot precisely localize specific nodes, they remain crucial indicators of long-running, persistent APT campaigns. This is because during an APT attack, adversaries may preserve node-level attributes, but they cannot maintain the benign program’s graph structure.
Finally, the interpretable surrogate DT enables the construction of actionable, security-aware explanations for individual predictions. Using ProvExplainer’s global surrogate DT and participating node’s system attributes (e.g., process names, file names, and IP/PORT), domain experts can leverage this enriched contextual information to understand the decision-making process of the GNN.
VI Evaluation
In this section, we evaluate ProvExplainer’s effectiveness in explaining stealthy attacks. We aim to answer the following research questions (RQs):
  • RQ1:
    Explanation Accuracy. Can ProvExplainer accurately explain APT and Fileless Malware detections (§ VI-D and § VI-E)?
  • RQ2:
    Comparison with SOTA GNN Explainers. How do ProvExplainer’s explanations compare against SOTA GNN explainers (e.g., GNNExplainer [31], PGExplainer [32], and SubgraphX [33]) (§ VI-F)?
  • RQ3:
    Explainability Effectiveness. How much actionable are ProvExplainer’s explanations (§ VI-G)?
VI-A Evaluation Protocols
We evaluate ProvExplainer on a different APT detection and Fileless Malware datasets: (1) publicly available APT attack simulations [darpa], (2) an APT dataset from a recent study [46], and (3) execution traces of Fileless Malware [47]. We implement two general-purpose SOTA GNN-based detector models, GAT [68] and GraphSAGE [69], following the methodology adopted in recent explanation literature [60, 62]. Our focus is to first thoroughly understand and explain IDS that directly leverage GNNs for security-critical tasks, before addressing more complex IDS architectures that use GNNs to extract graph embeddings for downstream anomaly detection, as discussed in [26]. Accordingly, we do not compare against recent GNN-based IDS [56, 20, 25, 55, 19, 21, 27, 26] that rely on custom graph embeddings and employ a separate model for final detection.
We conducted an ablation study to understand the impact of different feature sets (e.g., subgraph patterns and graph structural features). We also evaluate the explanations given by ProvExplainer against SOTA GNN explainers [31, 32, 33]. Finally, we conduct a non-trivial human-centered evaluation to validate the interpretability of the explanations from a security-analyst perspective. In § VII, we present case studies explaining APT detection and offer in-depth insights into how ProvExplainer provide security-relevant explanations.
VI-B Evaluation Metrics.
In our evaluation of ProvExplainer, we focus on three critical aspects. The first is the agreement of the surrogate DTs with the GNN model, which we measure using the weighted macro averaged (WMA) F1 score of the surrogate DT’s predictions with respect to the GNN’s predictions. The agreement metric gauges the faithfulness of the DT in replicating the conclusions of GNNs. WMA F1 score helps to account for the data imbalance issue present in the anomaly detection datasets.
[…]
To evaluate ProvExplainer’s and SOTA GNN explainers’ proficiency in identifying security-relevant entities, we define precision and recall metrics with respect to documented entities § II-D. Let $\mathcal{D}$ be the set of documented entities and using the above definitions for $\mathcal{V}_{i}$ and $k$. Precision is the proportion of explanation nodes that are documented, and recall is the fraction of documented entities retrieved: $precision(V_{k},k,\mathcal{D})=\frac{|V_{k}\cap\mathcal{D}|}{k}$, and $recall(V_{k},\mathcal{D})=\frac{|V_{k}\cap\mathcal{D}|}{|\mathcal{D}|}$.
Lastly, to assess the actionability of the explanations from a security analyst’s perspective, we introduce a graph traversal distance metric, where we compute the shortest-path distance between each node recommended by an explainer and its corresponding ground-truth node (§ II-D). Let $\mathcal{V}_{\text{GT}}$ denote the set of ground-truth nodes and $\mathcal{V}_{\text{expl}}$ the set of source nodes highlighted by the explainer. The distance metric is defined as: $distance_{traversal}=\frac{1}{|\mathcal{V}_{\text{GT}}|}\sum_{i=1}^{|\mathcal{V}_{\text{GT}}|}d(v^{\text{expl}}_{i},v^{\text{GT}}_{i})$, where $d(v^{\text{expl}}_{i},v^{\text{GT}}_{i})$ is the shortest-path node distance between the $i$-th recommended node and its matched ground-truth node in the graph. This node-based distance is then normalized by the number of ground-truth nodes $|\mathcal{V}_{\text{GT}}|$, ensuring a consistent scale across graphs. The intuition behind this metric is that a lower distance indicates the explainer’s recommendation is closer to the true malicious entities, thereby reducing the analyst’s cognitive and temporal burden during an investigation.
(a) DARPA APT Dataset [darpa]
(b) APT Dataset from [46]
(c) Fileless Malware from [47]
Figure 3: Agreement (higher is better) measured using WMA F1 of surrogate decision trees versus the GAT model across different feature sets.
VI-C Evaluation Tasks
APT Detection. We utilized the DARPA Transparent Computing (TC) Data Releases [darpa] and previous literature [46] for APT attack detection. This dataset, encompassing various OSes, provides a comprehensive basis for advanced security research. The DARPA APTs were designed to attack a system which consists of long-running processes and captured the stealthy attack vectors frequently employed by advanced adversaries. We particularly focused on three DARPA datasets used by previous studies [56, 19, 20, 21]: FiveDirections, Trace, and Theia. However, these tasks were conducted in a simulated environment and only lasted for two weeks, involving a limited number of hosts. Therefore, we also evaluated against the APT scenarios (e.g., Enterprise APT and Supply-Chain APT) conducted by [46], which were performed with realistic benign background workloads. Interested readers can find detailed statistics presented in Table IV.
Fileless Malware Detection. For Fileless Malware detection, we targeted a family of stealthy malware samples that impersonate benign programs, evading conventional security solutions but are detectable by GNN-based provenance analysis. The malware samples were chosen in accordance with guidelines from the literature [70, 71] to minimize experimental bias and ensure freshness. The Fileless Malware dataset includes various categories [70], including banking Trojans, spyware, ransomware, and malware installers; detailed statistics presented in Table IV in the appendix.
The Banking banking trojan [70] steals banking credentials from victim machines and spreads through spam and compromised download links. To camouflage interactions with .dlls and temporary files, it masquerades as svchost.exe. The Spyware, Ulise [70], is a multi-purpose Trojan that can establish remote access connections, capture keyboard input, collect system information, upload files, drop other malware into the infected system, and perform encryption. Because it propagates through the network sockets and interacts with many system files, it masquerades as python.exe as it needs a target that exhibits diverse behavioral patterns.
Dataset
GAT
Surrogate DT
(agree w/ GAT)
GraphSAGE
Surrogate DT
(agree w/ GraphSAGE)
DARPA APT Dataset [darpa]
FiveDirections 0.82 0.88 0.82 0.85
Trace 0.93 0.93 0.87 0.87
Theia 1.00 1.00 0.94 0.94
Average 0.92 0.93 0.88 0.89
APT Dataset from [46]
Enterprise 0.82 0.95 0.81 0.94
Supply-Chain 0.82 0.94 0.80 0.97
Average 0.82 0.95 0.81 0.95
Fileless Malware from [47]
wscript.exe 1.00 1.00 0.95 0.95
cscript.exe 1.00 1.00 1.00 1.00
reg.exe 1.00 1.00 1.00 1.00
python.exe 0.99 0.99 0.99 0.99
explorer.exe 0.98 0.98 1.00 0.98
netsh.exe 0.98 0.98 0.98 0.98
net.exe 0.97 0.97 0.97 0.97
rundll32.exe 0.96 0.96 0.96 0.96
mshta.exe 0.95 0.95 0.85 0.85
schtasks.exe 0.93 1.00 0.93 1.00
conhost.exe 0.89 0.89 0.93 0.93
svchost.exe 0.86 1.00 0.86 1.00
Average 0.96 0.98 0.95 0.97
Table I: Surrogate DTs have high agreement ($\geq$ 0.85) with the decisions of GNN measured using the WMA F1 score.
VI-D Graph Structural Feature Evaluation
To answer RQ1, Table I demonstrates the effectiveness of surrogate DTs in mirroring the decision process of GNN models like GAT and GraphSAGE . Agreement between the surrogate DT and the GNN is an important metric for two reasons: (1) surrogate explanations are only valid when the surrogate agrees with the GNN, and (2) high agreement indicates that the surrogate model is a good approximation of the GNN’s decision-making process.
ProvExplainer exhibits the highest agreement(>> 97%) with the GNN models on the Fileless Malware datasets, but the APT datasets only showed good agreement (>> 93%). The surrogate DTs’ superior performance in Fileless Malware datasets can be attributed to ProvExplainer’s ability to capture local subgraph patterns with high efficiency but APT attacks that are long running and utilize evasive tactics (noted by previous studies [46, 47]) are harder to detect and explain. The dataset such as FiveDirections and conhost.exe are dominated by attackers using stealthy techniques such as living-off-the-land, which involve memory object interactions that are currently not captured in our provenance graphs.
The absence of the distinguishing features or rare features that are exclusively present in the testing dataset impairs the effectiveness of ProvExplainer. We observed that GAT outperforms GraphSAGE in majority of the datasets. This trend is attributed to GAT ’s ability to assign varying importance to different neighborhood structures and capture long-range structural dependencies through its attention mechanism. Conversely, GraphSAGE , which samples a fixed-size neighborhood, overlooks subtle structural changes, such as those resulting from elusive APT attacks. ProvExplainer’s graph structural features enable surrogate DTs to efficiently approximate GNN models’ decision-making process on in-distribution data. Although the current features are sensitive to attacker’s TTPs, ProvExplainer is expandable to incorporate new features to extend support for the new TTPs.
Figure 4: Comparison of GAT explainers in identifying important structures using $\text{fidelity}^{+}$ (higher is better) and $\text{fidelity}^{-}$ (lower is better).
Figure 5: Effectiveness of GNN explainers at identifying documented entities, measured using precision and recall (higher is better for both).
VI-E Ablation Study
Figure 3 shows the contributions of different features to the overall surrogate DTs agreement of approximating a GAT model. Interested readers may find this study with GraphSAGE in the appendix, Figure 8. We separate the features into two groups: subgraphs and graph structural features. We evaluate each group individually and then we look at the their combined effect. We further separate the group into individual features (e.g., top-$K$ subgraphs) and look at the effect of each constituting features in the group. Furthermore, to explore the performance of different groups, we evaluate different graph structural features individually and collectively.
In the APT dataset, the top-$K$ subgraphs individually show different levels of agreement for both the Enterprise and Supply-Chain subsets, indicating that distinct subgraphs capture complementary signals. A similar trend is observed for the DARPA TC and Fileless Malware datasets, although the increase in agreement levels plateaus as $K$ approaches 4. This suggests that the highest-ranked subgraphs encode the most informative patterns, supporting our claim that the decision path extracted from the surrogate DT effectively captures the discriminative power of the GNN.
We examine individual structural features, certain metrics (i.e., eigenvector centrality and clustering coefficient) consistently exhibit high agreement in the APT datasets, particularly in DARPA TC (Trace and Theia) dataset. In contrast, other metrics (i.e., closeness and betweenness centrality) show low agreement across most APT datasets. This behavior arises because, in the context of APT attacks, centrality measures often encode information that misleads the model: benign graphs contain values similar to those of attack graphs. As attackers commonly leverage system processes [darpa] to do their bidding and system processes exhibit similar structural characteristics, which reduces their discriminative power.
In the APT dataset, where attackers employ stealthy techniques, which suggests the necessity of integrating subgraph structures with graph structural features to construct a comprehensive feature set as combined features provide stronger alignment than relying on individual features. ProvExplainer achieves the best performance across all datasets when all features are combined.
An interesting pattern is that the Fileless Malware dataset consistently performs exceptionally well, particularly with individual features like eigenvector centrality and clustering triangles, which achieve near-perfect alignment scores. This suggests that the structural properties in the Fileless Malware dataset are distinctively different compared to the benign programs they are masquerading.
VI-F ProvExplainer vs. SOTA Explainers
To answer RQ2, Figure 4 compares the $\text{fidelity}^{+}$ $\text{fidelity}^{-}$ and Figure 5 compares the precision/recall of the explanations derived from ProvExplainer and SOTA methods (GNNExplainer [gnnexplainer], PGExplainer [pgexplainer], and SubgraphX [subgraphx]). We analyze specific case studies in § VII.
$\text{Fidelity}^{+}$ and $\text{Fidelity}^{-}$. In the APT dataset from [46], ProvExplainer achieves the best performance for both $\text{fidelity}^{+}$ and $\text{fidelity}^{-}$ as seen in Figure 4. ProvExplainer and SubgraphX show competitive $\text{fidelity}^{+}$ results, outperforming both GNNExplainer and PGExplainer. For $\text{fidelity}^{-}$ , GNNExplainer performs the worst, as it fails to identify all the critical subgraph structures. Therefore, when graph structures deemed unimportant are removed, the predictions still change, resulting in high $\text{fidelity}^{-}$ value. ProvExplainer’s superior performance in the DARPA and APT datasets is attributed to its security-aware features, which provide a significant advantage in identifying and extracting critical nodes from provenance graphs. These nodes, when removed, lead to prediction changes, demonstrating ProvExplainer’s ability to capture essential graph components.
Additionally, ProvExplainer achieves high coverage of important nodes, reflected in its low $\text{fidelity}^{-}$ values, removing unimportant nodes does not affect predictions. In the Fileless Malware dataset, ProvExplainer demonstrates moderate performance at lower $k$ values (e.g., the total number of explanation nodes considered). However, as $k$ increases, ProvExplainer becomes the best performer in $\text{fidelity}^{+}$ , with SubgraphX ranking second. These results highlight that SubgraphX focuses on a limited number of structures and fails to provide comprehensive coverage of all the critical substructures essential for accurate predictions.
ProvExplainer’s $\text{fidelity}^{-}$ performance is hindered in the Fileless Malware dataset due to the unique substructures created by Fileless Malware attacks, which are not adequately captured by its security-aware features. If distinguishing subgraphs present in the test set are absent from the training set, these critical structures are missed, leading to gaps in the explanations. However, the $\text{fidelity}^{-}$ performance of other explainers declines more rapidly as $k$ increases in the Fileless Malware dataset. This suggests that general-purpose GNN explainers, which operate by calculating masks around edges and subgraphs, perform better when unique subgraph structures are present for each specific attack. It is important to note that ProvExplainer’s primary objective is not to serve as a zero-day vulnerability detector but rather to produce human-interpretable explanations.
Precision and Recall. In the APT datasets, ProvExplainer demonstrates superior precision and recall performance for both GAT and GraphSAGE networks compared to other GNN explainers as seen in Figure 5. This highlights ProvExplainer’s reliability in generating accurate explanations for the APT datasets by identifying relevant subgraph structures that include nodes mentioned in the ground truth. A similar trend is observed in Fileless Malware when explaining GAT , although ProvExplainer achieves the second-best precision and recall in this case. These results emphasize ProvExplainer’s strength in explaining complex APT datasets, where common TTPs from the MITRE framework are frequently reused (i.e., similar subgraphs are formed). However, ProvExplainer struggles with the Fileless Malware dataset, where unique structures in the testing set are absent in the training set, illustrating the impact of novel TTPs on explainer performance.
GNNExplainer performs well in the Fileless Malware dataset category due to its fidelity-driven approach, where GNNExplainer calculates masks for substructures for every instance, enabling it to effectively identify diverse structures. Similarly, SubgraphX shows improved precision and recall when explaining GraphSAGE compared to GAT , as SubgraphX focuses on identifying important subgraphs based on fixed-size node neighborhoods. These neighborhoods influence the message-passing algorithm of GraphSAGE , ensuring that the explanations generated by SubgraphX align closely with GraphSAGE ’s computational behavior.
ProvExplainer’s performance demonstrates its strength in the APT and Fileless Malware datasets, achieving both high precision and perfect recall. This suggests a strong alignment with the inherent characteristics and structure of these datasets. However, in the more complex APT dataset [46], all explainers, including ProvExplainer, face challenges in maintaining high precision. This limitation arises due to the dataset’s complexity and the novel TTPs employed by APT attackers during certain stages. These TTPs are represented in only a subset of the graphs. Consequently, when these specific graphs appear exclusively in the testing set, ProvExplainer lacks the corresponding subgraphs for explanation, resulting in missed patterns and a subsequent decline in precision.
Table II: Explanation actionability comparison between ProvExplainer and SOTA GNN explainers measured by human actionability metric. Absolute human actionability metric (top) and relative factor versus ProvExplainer (italic, bottom). Lower values reflect better actionable.
Dataset
GNNExplainer PGExplainer SubgraphX ProvExplainer
DARPA APT
Dataset [darpa]
3.29
[…]
2.60
[…]
2.36
[…]
2.18
[…]
APT Dataset
from [46]
1.97
[…]
1.71
[…]
1.85
[…]
1.24
[…]
Fileless Malware
from [47]
2.04
[…]
1.74
[…]
1.77
[…]
1.34
[…]
VI-G Human Actionability Study
[…]
VII Case Studies
To demonstrate ProvExplainer in a realistic setting, we analyze case studies from the DARPA [darpa] datasets. In each study, we use an explanation size of 40 nodes and refer to the GAT model. We list the most important features from the surrogate DT and qualitatively analyze the explanations from ProvExplainer and the SOTA GNN explainers. One more case study (§ A-E) and detailed analyses in the appendix (§ A-F).
VII-A FiveDirections: Browser Extension [darpa]
(a) Attack Graph.
(b) SOTA GNN explainations.
(c) Pattern #1.
Figure 6: FiveDirections: the attacker exploits the target via a malicious Firefox extension.
Description. An attacker targets the Firefox browser using a malicious extension to deploy the drakon malware. The attacker writes drakon directly to disk and exploits a compromised browser extension masquerading as a password manager to execute malicious powershell code, gaining deeper access and control over the system (Figure 6).
Features: Degree centrality and pattern #1.
System Interpretation. The pattern shown in 6(c) provides a detailed representation of the initial stages (i.e., initial compromise and privilege escalation) of APT attack lifecycle orchestrated by the Drakon APT, showcasing a sequence of file and process interactions. Process P1 (pass_mgr.exe) serves as the initial pivot point, leveraging File F1 (firefox.exe) to gain access to the system through a compromised browser backdoor. This initial stage involves P1 writing to multiple files, including File F2 (ntdll.dll), File F3 (sysmain.sdb), File F4 (passwordfile.dat), and File F5 (RICHED20.DLL). These operations signify the deployment of malicious shellcode, preparation for privilege escalation, and establishment of a foothold within the system.
The transition to Stage 2 is marked by the activities of Process P2 (pass_mgr.exe), which reads from File F5 (RICHED20.DLL) and writes to critical files such as File F3 and File F4. These interactions indicate the progression to privilege escalation and credential harvesting, leveraging tools like Copykatz and Drakon APT’s elevate driver. The repeated use of the same process executable (pass_mgr.exe) across different stages points to sophisticated APT TTPs such as process injection or executable reuse.
Since the attack graph contains high quantities of file read and write flowers for a relatively small number of nodes between these flowers, it leads to a high degree of centrality, which means a high amount of process interaction between static artifacts like files. The graph’s structural measures complement the subgraph’s interpretation and identifies the TTPs employed by Drakon APT.
ProvExplainer vs. SOTA explainers. 6(b) compares the explanations of ProvExplainer and those of SOTA GNN explainers, focusing on their efficacy in identifying security-aware elements. ProvExplainer excels by highlighting the malware replicating itself from template files and accessing sensitive system files. ProvExplainer isolates security-relevant structures in the graph, significantly increasing the end-user trust in the detection.
GNNExplainer identifies the malware template file and the malicious extension. This effectiveness stems from GNNExplainer’s method of searching for important edges. When this GNNExplainer isolates the pivot structure, the graph becomes disjoint, leading to a change in prediction. This results in a high information gain, the core metric for GNNExplainer. On the other hand, PGExplainer and SubgraphX reveal commonalities across attack graphs, such as the identification of key system library accesses required for malware operation. SubgraphX’s effectiveness varies due to its Monte Carlo Tree Search (MCTS), suggesting benefits to incorporating domain-specific insights into SubgraphX’s scoring function.
VIII Discussion and Future Work
Limitations. ProvExplainer depends on previously observed graph patterns, as the surrogate DT is trained on discriminant subgraph patterns mined from an existing dataset. Therefore, ProvExplainer is unlikely to perform well on attacks that contain previously unseen subgraph patterns. We recognize the inherent limitations of our approach and, therefore, set explicit boundaries for ProvExplainer’s interpretive range, aiming to validate detection results and provide actionable intelligence for known attacks. Preconditions that may prevent ProvExplainer from explaining some attacks: (1) the detection model must detect the attack, (2) the attack graph contains distinct subgraphs identified during the mining phase, and (3) the attack’s distinctive subgraphs are small enough to be feasibly pattern-matched.
Adversarial Manipulation of Graph Features. The explainer models used in ProvExplainer is separate from the black-box detection model. Fooling both the surrogates and the detection model is the ideal case for an attacker seeking to avoid suspicion, and an attacker who can attack the explainer this way can also attack the detection model directly [goyal2023sometimes, 46]. Creating robust explanation framework that can withstand adversarial manipulation is a critical open research problem, and is beyond the scope of this work.
IX Conclusion
We introduced ProvExplainer, a framework designed to address a critical gap in explainable ML for cybersecurity. It provides transparent, interpretable insights into GNN-based IDS in the system provenance domain. By leveraging security-aware discriminative subgraph patterns and graph structural features, ProvExplainer facilitates instance-level explanations that consider both local and global graph properties. This dual-level interpretability is essential for removing the opacity and help with the adoption of GNN-based IDS. ProvExplainer is a critical advancement in explainable AI for the security domain and a scalable framework for real-world deployment.
Appendix A Appendix
Table III: Node and edge types for in-house system provenance with associated attributes.
Types Attributes
Nodes
process signature, executable name, pid
file owner (uid, gid), name, inode
socket (IP) dstip, srcip, dstport, srcport, type
Edges
process $\rightarrow$ process command args, starttime
process $\rightarrow$ file read, write, amount
process $\rightarrow$ IP address send, recv, amount
A-A Implementation
Our data collection module can operate on Windows systems by using the Windows ETW [72] and on Linux system by using Linux auditd [73] frameworks to collect relevant system calls regarding files, processes and network sockets. These include system calls for (1) file operations (e.g., read(), write(), unlink()), (3) network socket operations (e.g., connect(), accept()), (4) process operations (e.g., create(), exec(), and exit()). The system-level data is stored in a PostgresSQL database. Table III summarizes the event collection schema for our in-house deployment. While it includes major nodes and event types annotated with textual and numerical attributes, detailed entries of the event collection differ across OS environment.
ProvExplainer is implemented using python and graph generation framework is implemented in java. The GNN training and evaluation pipelines are implemented using Deep Graph Library (DGL) [58] framework, and the surrogate DTs are implemented using sklearn [74]. We used the frequent subgraph miner, GERM [39], to mine patterns with a minimum support of 1 and a maximum subgraph size of 10. The discriminant subgraph patterns were selected by maximizing the scoring function. The scoring function’s hyperparameters and DT were empirically tuned using grid search by iteratively evaluating ProvExplainer’s performance using Fidelity+/- and Precision/Recall metric on cross-validation folds.
A-B Dataset Statistics
Table IV: APT and Fileless Malware graph statistics.
Applications
# of Benign
Graphs
# of Anomaly
Graphs
Avg # of Benign
Nodes / Edges
Avg # of Anomaly
Nodes / Edges
DARPA APT Dataset [darpa]
Trace 1883 8 735.35 / 957.56 836.15 / 946.75
Theia 2858 9 559.47 / 979.59 913.91 / 987.31
FiveDirections 954 13 906.22 / 971.91 959.43 / 973.63
Average 1898.33 10.00 669.81 / 982.18 777.08 / 1011.19
APT Dataset from [46]
Enterprise 3079 1836 90.22 / 85.13 73.73 / 76.88
Supply-Chain 3212 1092 65.02 / 40.77 61.09 / 54.33
Average 3145.50 1464.00 55.94 / 83.29 58.38 / 51.69
Fileless Malware from [47]
wscript.exe 399 40 66.18 / 59.94 44.10 / 56.45
cscript.exe 876 11 88.56 / 81.58 87.36 / 101.54
reg.exe 309 116 60.18 / 52.93 78.91 / 131.87
python.exe 15585 426 89.95 / 83.15 57.98 / 79.33
explorer.exe 399 40 66.18 / 59.94 44.10 / 56.45
netsh.exe 621 7 44.08 / 37.49 38.42 / 52.28
net.exe 621 7 44.08 / 37.49 38.42 / 52.28
rundll32.exe 1632 443 52.42 / 46.42 51.97 / 81.31
mshta.exe 876 11 88.56 / 81.58 87.36 / 101.54
schtasks.exe 399 40 66.18 / 59.94 44.10 / 56.45
conhost.exe 309 116 60.18 / 52.93 78.91 / 131.87
svchost.exe 1632 443 52.42 / 46.42 51.97 / 81.31
Average 2667.00 90.84 60.44 / 54.15 63.72 / 89.93
Anomaly Detection Dataset. The dataset statistics for the anomaly detection dataset is seen in Table IV. The Fileless Malware samples were downloaded from [30] and selected from recent studies, [75, 70, 71, 47]. The benign provenance graphs for the anomaly detection dataset were sourced from the DARPA dataset, which contained 5,695 benign graphs with an average of 669.81 vertices and 982.18 edges (Table IV); the Fileless Malware dataset, which contained 19,422 benign graphs with an average of 60.44 vertices and 54.15 edges; and the APT dataset, which contained 6,291 benign graphs with an average of 55.94 vertices and 83.29 edges. The corresponding anomaly graphs are sourced from the same datasets: The DARPA dataset contained 30 anomalous graphs with an average of 777.08 vertices and 1011.19 edges (Table IV); the Fileless Malware contained 1,043 anomalous graphs with an average of 63.72 vertices and 89.93 edges; and the APT dataset contained 2,928 graphs with an average of 58.38 vertices and 51.69 edges.
A-C Ablation Study (using GraphSAGE)
(a) DARPA APT Dataset [darpa]
(b) APT Dataset from [46]
(c) Fileless Malware from [47]
Figure 7: Agreement (higher is better) measured using WMA F1 of surrogate decision trees versus the GraphSAGE model across different feature sets.
The ablation study utilizing the GraphSAGE network revealed patterns consistent (as shown in Figure 7) with those observed in the ablation study using GAT network § VI-E. It was again noted that for Fileless Malware datasets the graph structural features provided better feature selection than subgraph patterns. This pattern is observed because the malware graphs are distinctively different when compared against their benign counter part. Finally, the pattern that subgraph patterns with the graph structures provided the best feature selection.
A-D ProvExplainer vs. SOTA Explainers
Figure 8: Effectiveness of GraphSAGE explainers in identifying prediction-relevant graph structures using $\text{fidelity}^{+}$ and $\text{fidelity}^{-}$ .
Similar to the GAT -based results, ProvExplainer demonstrates superior performance across various datasets. The complexity of the provenance graphs, defined by their size and the nature of resource interactions, significantly influences the effectiveness of different explainers on specific datasets. For instance, on provenance graphs such as the APT graphs from [46], PGExplainer achieves the second-best performance. However, for the more complex or noisy DARPA APT graphs, GNNExplainer emerges as the second most effective method.
A-E Case Study: FiveDirections (Copykatz) [darpa]
(a) Attack Graph.
(b) SOTA GNN explainations.
(c) Pattern #2.
Figure 9: FiveDirections: the attacker gains C2 connections and installs Copykatz through a Firefox exploit.
Description. In a sophisticated attack, a hijacked version of usdoj.gov exploits Firefox to deploy drakon malware to the victim host. Then, drakon uses the elevate driver to escalate privileges and masquerade as the runtimebroker.exe system program. Finally, the malicious runtimebroker.exe instance connects to a command and control (C2) server to download and execute Copykatz (an older version of Mimikatz) to harvest and exfiltrate host credentials, as illustrated in Figure 9.
Features: Degree centrality and pattern #2.
System Interpretation. The pattern shown in 9(c) highlights a multi-stage attack sequence characteristic of the Firefox Drakon APT focusing on elevating the privilege using Copykatz. Process P3 (firefox.exe) initiates the attack by exploiting its backdoor to download the payload, File F1 (ctfhost2.exe) that contains the drakon malware. This payload is subsequently executed by Process P1 (runtimebroker.exe), which modifies critical system libraries, File F2 (ntdll.dll) and File F3 (kernel32.dll). These modifications indicate the preparation of the system environment for privilege escalation and sustained malicious operations. Process P2 (svchost.exe) interaction with File F2 emphasizes its role in leveraging system-level components to propagate the attack, consistent with the exploitation of driver signing for privilege elevation.
The graph structure captures the interdependencies between processes and system-critical files, illustrating the deliberate flow of the attack. By chaining interactions, such as runtimebroker.exe modifying ntdll.dll and svchost.exe subsequently accessing it, the graph provides insights into how the malware manipulates system operations. Furthermore, the reuse of Firefox as both an initial access vector and a conduit for kernel32.dll interactions highlights the advanced tactics employed to achieve persistence and exploitation. These observations align closely with the described TTPs of the Firefox Drakon Copykatz APT, offering a clear mapping of graph elements to specific stages of the attack lifecycle. The attack graph contains high quantities of file read and writes flowers for a relatively small number of process nodes acting as bridges between the flowers, so it leads to a high degree of centrality.
ProvExplainer vs. SOTA explainers. GNNExplainer and PGExplainer pinpoint the stage where multiple Firefox processes connect to the C2 servers. This activity traces back to the malicious runtimebroker instance. Notably, SubgraphX detects an alternate trajectory where a Firefox process, instead of reaching out to external servers, spawns another process aimed at local content manipulation, showcasing the malware’s versatility in engaging with both external and internal resources for its objectives. While SOTA explainers have demonstrated proficiency in identifying the final stages of this data breach, only ProvExplainer effectively captured both the initial infection and its privilege escalation attempts. This distinction highlights the importance of recognizing early-stage indicators for root cause analysis.
A-F In-Depth Analysis of Case Studies
Trace: Phishing E-mail [darpa]. The pattern #2 captured the staging behavior of the Trojan. The Trojan was downloaded from www.nasa.ng, executed and replicated itself within the system. Specifically, the malware named nasa.ng is placed in /home/admin/.mozilla/firefox/ and /usr/local/firefox-54.0.1/obj-x86_64-pc-linux-gnu/. The Trojan created new malware with benign names such as firefox to effectively evade detectors, to replicate unhindered and overloaded the system with malicious processes. After the malware successfully staged, the malware masquerading as /bin/sh to read the malicious script staged in (/etc/update-motd.d/00-header/) and executed it to create multiple copies of itself. The malware reads various system configuration files present in /etc/protocols/, /etc/lsb-release/, and /etc/hosts.deny/. Reading sensitive system configuration files are essential to build the system profile. The pattern #5 captured the dependency correlation of the malware created that inherited configurations.
[…]
GNNExplainer was able to correctly capture the staging behavior where the malware from nasa.ng read shared library (/usr/lib/x86_64-*/*.so.*) and cache file (/usr/share/applications/mimeinfo.cache, /usr/lib/x86_64-*/*/loaders.cache). PGExplainer incorrectly indicated benign substructures, but SubgraphX correctly captured the inheritance behavior of firefox.exe executing multiple times with the argument file http://www.nasa.ng/, to create the malware clones from the template. SOTA explainers missed the malware’s ultimate goal of reading sensitive files, which was only captured by ProvExplainer.
FiveDirections: Browser Extension [darpa]. In the context of the described attack, where the drakon malware exploits the firefox.exe browser through a rogue extension (pass_mgr) the pattern #1 is created. Multiple instances of the malware process are created, each reading from the malicious files: passwordfile.dat and pass_mgr.exe. Additionally, they access essential dictionary files (en-US.aff) and cryptographic libraries (bcryptprimitives.dll ) to maintain operational consistency, allowing them to conduct their malicious activities efficiently. The pattern #2 captures the data extraction behavior through sensitive file accesses. These malware access sensitive information and system configuration files like WindowsShell.Manifest, and wintrust.dll, along with initial malware files (addons, tzres.dll, userenv.dll).
[…]
FiveDirections: Copykatz [darpa]. The pattern #2 starts its kill chain stage of initial access [76] by writing and executes a malware masquerading as firefox.exe which contains the Mimikatz and Copykatz modules. ntdll.dll [77] includes multiple kernel-mode functions which enables the “Windows Application Programming Interface (API)” and bcryptprimitives.dll [78] contain functions implementing cryptographic primitives, which are essential for Mimikatz [40] and Copykatz.
The firefox.exe malware write its payload into temporary files, such as ctfhost.exe, virtuous and tropical. These files are then executed by malware in the system creating pattern #5. Subsequently, another malicious instance of the firefox.exe malware initiates a chain or read-execute cycle by executing the malware template as well as making external IP connections. This dependency correlation between the malware and its parent is characterized by inheritance, where the children require the same system library files as the parent to function correctly. The DLLs involved are read from ProgramFiles and System32.
GNNExplainer and PGExplainer correctly identified the data exfiltration stage where four malicious firefox.exe make C2 connections to the external IPs (202.179.137.58 and 217.160.205.44). firefox.exe was invoked by malware masquerading as a popular system program, runtimebroker.exe. GNNExplainer correctly masks out the process creation edges of the children firefox.exe, to create disjoint graphs and leading to prediction change. SubgraphX also highlighted firefox.exe processes that are created by runtimebroker.exe, but instead of making external connection it creates other firefox.exe for rendering content from localhost (127.0.0.1).